# Credal Learning Theory

### v3

@@ 5 Conclusions
In this paper, we laid the foundations of a more general statistical learning theory, that we called Credal Learning Theory (CLT). We generalized some of the most important results of classical Statistical Learning Theory (SLT) to allow for drift and misspecification of the data generating process. We did so by considering sets of probabilities (credal sets), instead of single distributions. The modeling effort needed to elicit credal sets is paid off in terms of the tightness of the resulting bounds. This is just the first step in an ongoing endeavor to develop a comprehensive apparatus that extends all the classical SLT results to a credal sets system. In the future, we plan to further our undertaking, for instance by (i) modeling the epistemic uncertainty induced by domain variation through random sets rather than credal sets, (ii) comparing our method with robust learning (Cianfarani et al., 2022), and (iii) extending our results to different losses. In this work, we focused solely on the zero-one loss to be able to directly build on the classical results in Liang (2016, Chapter 3). We also intend to validate empirically our findings. For now, a synthetic experiment showing that the bound in Theorem 4.1 is tighter than the classical SLT one, can be found in Appendix B.
@@ Appendix A Proofs
The proof builds on that of Liang [2016, Theorem 7]. Fix any $\epsilon>0$, and any $P\in\mathcal{P}$. Assume that the training dataset is given by $n$ i.i.d. draws from $P$. By Liang [2016, Equations (158) and (186)], we have that
$\begin{split}\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}&\leq\mathbb{P}\left[\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]\\ &<|\mathcal{H}|\cdot 2\exp\left(-2n\left(\frac{\epsilon}{2}\right)^{2}\right)\\ &\doteq\delta(\epsilon).\end{split}$ (11)
Notice though, that we can improve on this bound, since we know that $P\in\mathcal{P}$, a credal set. Let $B^{\prime}_{P}\doteq\{h\in\mathcal{H}:|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}$ be the set of “bad hypotheses” according to $P$. Then, it is immediate to see that
$\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{P}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|.$
Notice though that we do not know $P$; we only know it belongs to $\mathcal{P}$. Hence, we need to consider the set $B^{\prime}_{\mathcal{P}}$ of bad hypotheses according to all the elements of $\mathcal{P}$, that is, $B^{\prime}_{\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P\in\mathcal{P}\text{,}|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P\in\mathcal{P}}B^{\prime}_{P}$. Since $\mathcal{P}$ is a credal set, by the Bauer Maximum Principle and the linearity of the expectation operator we have that $B^{\prime}_{\mathcal{P}}=B^{\prime}_{\text{ex}\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P^{\text{ex}}\in\text{ex}\mathcal{P}\text{, }|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B^{\prime}_{P^{\text{ex}}}$. Hence, we obtain
$\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|.$
@@
$\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}$ $\leq\mathbb{P}\left[\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]$
@@
Fix any $\delta\in(0,1)$. In Liang [2016, Theorem 9], the author shows that for a fixed probability measure $P$ on $\mathcal{X}\times\mathcal{Y}$, we have that
@@
Appendix B A Synthetic Experiment on Theorem 4.1
In this section, we perform a synthetic experiment to show that the bound we find in Theorem 4.1 is indeed tighter than the classical SLT one reported in Corollary 4.3. In recent literature, studies by Kacham and Woodruff [2022], Amit et al. [2022], Li and Liu [2021] have conducted synthetic experiments in a similar manner. These works are mainly theoretical in nature, but they also acknowledge the importance of experimental validation with preliminary analysis.
Experiment 1: Let the available training sets be $D_{1},D_{2},D_{3}$. Assume, for simplicity, that $\Omega=\mathcal{X}\times\mathcal{Y}=\{x\}\times\mathbb{R}\simeq\mathbb{R}$. Suppose that we specified the likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$. Call $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ their respective probability measures, and derive the credal set $\mathcal{P}$ as we did in footnote 7. That is, let $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{3})$. We determine the credal set in this way because it is then easy to find its extreme elements $\text{ex}\mathcal{P}$. Indeed, it is immediate to notice that $\text{ex}\mathcal{P}=\{\mathcal{L}_{i}\}_{i=1}^{3}$. Let now $D_{N+1}\equiv D_{4}$ be a collection of $n$ samples from $P^{\text{true}}\equiv\mathcal{L}_{2}\in\mathcal{P}$. The hypothesis space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and we consider the zero-one loss $l$ as we did in the main portion of the paper.
@@
As we can see from Table 1, our bound $\epsilon^{\star}(\delta)$ improves on the classical SLT one $\epsilon_{\text{UB}}(\delta)$. Table 1 also tells us that bound $\epsilon^{\star}(\delta)$ is tighter than $\epsilon_{\text{UB}}(\delta)$ when the sample size $n=|D_{4}|$ is small, and then $\epsilon^{\star}(\delta)$ becomes progressively closer to $\epsilon_{\text{UB}}(\delta)$ as $n=|D_{4}|$ increases. This same pattern is observed when the extrema of the credal set are closer to each other. Indeed, in Table 2 we repeat the experiment and choose as extrema of $\mathcal{P}$ three measures whose pdf’s are three Normals $\mathcal{N}(-0.1,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(0.1,1)$. [^12] Of course, they are much closer to each other than the Normals $\mathcal{N}(-5,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(5,1)$, e.g. in the Total Variation metric. The reason for this behavior is the following. With few available samples, that is, when $n=|D_{4}|$ is low, Credal Learning Theory is able to leverage the evidence encoded in the credal set, and hence to derive a tighter bound than classical Statistical Learning Theory. When the sample size is large, that is, when $n=|D_{4}|$ is high, the classical bound $\epsilon_{\text{UB}}(\delta)$ itself is very small. This because the amount of evidence available is large, and so $1/n\sum_{i=1}^{n}l((x_{i},y_{i}),h)$ well approximates $\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\text{true}}(\text{d}(x,y))$. In turn, since $\epsilon_{\text{UB}}(\delta)$ is already very small, then the CLT bound $\epsilon^{\star}(\delta)$ we derive cannot improve greatly on it. Hence, their values are close together, despite $\epsilon^{\star}(\delta)$ being slightly tighter. The code for this experiment is available upon request.
Table 1: Results of experimental evaluation of our bound tightness. Here the hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$.
@@
Table 2: Results of experimental evaluation of our bound tightness. Here the hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-0.1,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(0.1,1)$.
@@
For the experiment, we selected a standard Gaussian distribution $\mathcal{N}(0,1)$ (mean 0, standard deviation = 1) to generate data. Similarly to Experiment 1, (i) the hypothesis space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and (ii) we assume a zero-one loss function. The latter is used to evaluate the performance of the classifiers. For each run, we generated a training set and a test set from the standard Gaussian distribution. At training time, for each hypothesis $h$ in $\mathcal{H}$, we calculate the empirical risk on the training set using the zero-one loss and identify the hypothesis $\hat{h}$ that minimizes such risk (the empirical risk minimizer). At test time, we compute the empirical risk $L_{P}(\hat{h})$ of $\hat{h}$ on the test set as shown in Table 3.131313Here $P$ denotes the probability measure whose pdf is the standard Normal density.
We calculate the upper bound $\epsilon_{\text{UB}}(\delta)$ on the empirical risk based on Corollary 4.3, which is a function of the number of hypotheses in $\mathcal{H}$, the value of $\delta$, and the number $n$ of training samples. This is the classic SLT bound. The experiment is run $1000$ times with specified numbers of training and test samples, and a $\delta$ value of $0.05$ to check whether the condition (empirical risk on the test set is upper bounded by the theoretical bound) is satisfied in any of the $1000$ trials. The experimental results in Table 3 validate the classical SLT bound by repeatedly testing it on randomly generated data. The code for this experiment is also available upon request.
Table 3: Results of experimental evaluation. The hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$.
@@
Then, suppose that $\epsilon_{1}=0.2$, $\epsilon_{2}=0.3$, $\epsilon_{3}=0.1$, and $\epsilon_{4}=0.25$, so that $\mathscr{L}_{1}=\{P:P=0.8\mathcal{L}_{1}+0.2Q\text{, }\forall Q\}$, $\mathscr{L}_{2}=\{P:P=0.7\mathcal{L}_{2}+0.3Q\text{, }\forall Q\}$, $\mathscr{L}_{3}=\{P:P=0.9\mathcal{L}_{3}+0.1Q\text{, }\forall Q\}$, and $\mathscr{L}_{4}=\{P:P=0.75\mathcal{L}_{1}+0.25Q\text{, }\forall Q\}$. By Wasserman and Kadane [1990, Example 3], we know that, if we put $\mathcal{P}=\text{Conv}(\cup_{i=1}^{4}\mathscr{L}_{i})$, the following holds
$\underline{P}(A)=\begin{cases}\min_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A),&\forall A\neq\Omega\\ 1,&\text{if }A=\Omega\end{cases}\quad\text{and}\quad\overline{P}(A)=\begin{cases}\max_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)+\epsilon_{i},&\forall A\neq\emptyset\\ 0&\text{if }A=\emptyset\end{cases}.$

### v4

@@ 5 Conclusions
In this paper, we laid the foundations of a more general Statistical Learning Theory (SLT), that we called Credal Learning Theory (CLT). We generalized some of the most important results of classical SLT to allow for drift and misspecification of the data-generating process. We did so by considering sets of probabilities (credal sets), instead of single distributions. The modeling effort needed to elicit credal sets is paid off in terms of the tightness of the resulting bounds.
Limitations. (i) We only consider the zero-one loss in our results (we did so to be able to directly build on the classical results in Liang [44, Chapter 3]). (ii) We assume that the true distribution which the elements of the new test set $D_{N+1}$ are sampled from, belongs to the credal set that we derive at training time.
Future work. In the future, we plan to further our undertaking, for instance by (i) modeling the epistemic uncertainty induced by domain variation through random sets rather than credal sets, (ii) comparing our method with robust learning [23], (iii) extending our results to different losses, and (iv) deriving PAC-like guarantees on the correct distribution $P$ being an element of the credal set $\mathcal{P}$. We also intend to validate our findings on real datasets.
@@ Appendix A Proofs
The proof builds on that of Liang [44, Theorem 7]. Fix any $\epsilon>0$, and any $P\in\mathcal{P}$. Assume that the training dataset is given by $n$ i.i.d. draws from $P$. By Liang [44, Equations (158) and (186)], we have that
$\begin{split}\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}&\leq\mathbb{P}\left[\sup_{h\in\mathcal{H}}\left|\hat{L}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]\\ &<|\mathcal{H}|\cdot 2\exp\left(-2n\left(\frac{\epsilon}{2}\right)^{2}\right)\\ &\doteq\delta(\epsilon).\end{split}$ (8)
Notice though, that we can improve on this bound, since we know that $P\in\mathcal{P}$, a credal set. Let $B^{\prime}_{P}\doteq\{h\in\mathcal{H}:|\hat{L}({h})-L_{P}(h)|>\epsilon/2\}$ be the set of “bad hypotheses” according to $P$. Then, it is immediate to see that
$\sup_{h\in\mathcal{H}}\left|\hat{L}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{P}}\left|\hat{L}({h})-L_{P}(h)\right|.$
Notice though that we do not know $P$; we only know it belongs to $\mathcal{P}$. Hence, we need to consider the set $B^{\prime}_{\mathcal{P}}$ of bad hypotheses according to all the elements of $\mathcal{P}$, that is, $B^{\prime}_{\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P\in\mathcal{P}\text{,}|\hat{L}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P\in\mathcal{P}}B^{\prime}_{P}$. Since $\mathcal{P}$ is a credal set, by the Bauer Maximum Principle and the linearity of the expectation operator we have that $B^{\prime}_{\mathcal{P}}=B^{\prime}_{\text{ex}\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P^{\text{ex}}\in\text{ex}\mathcal{P}\text{, }|\hat{L}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B^{\prime}_{P^{\text{ex}}}$. Hence, we obtain
$\sup_{h\in\mathcal{H}}\left|\hat{L}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}({h})-L_{P}(h)\right|.$
@@
$\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}$ $\leq\mathbb{P}\left[\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]$
@@
Fix any $\delta\in(0,1)$. In Liang [44, Theorem 9], the author shows that for a fixed probability measure $P$ on $\mathcal{X}\times\mathcal{Y}$, we have that
@@
Appendix B Synthetic Experiments on Theorems 4.1 and 4.5
In this section, we perform synthetic experiments to show that the bounds we find in Theorems 4.1 and 4.5 are indeed tighter than the classical SLT ones reported in Corollaries 4.3 and 4.7, respectively. In recent literature, studies by Amit et al. [7], Kacham and Woodruff [38], Li and Liu [43] have conducted synthetic experiments in a similar manner. These works are mainly theoretical in nature, but they also acknowledge the importance of experimental validation with preliminary analysis.
Experiment 1: Let the available training sets be $D_{1},D_{2},D_{3}$. Assume, for simplicity, that $\Omega=\mathcal{X}\times\mathcal{Y}=\{x\}\times\mathbb{R}\simeq\mathbb{R}$. Suppose that we specified the likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$. Call $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ their respective probability measures, and derive the credal set $\mathcal{P}$ as we did in footnote 7. That is, let $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{3})$. We determine the credal set in this way because it is then easy to find its extreme elements $\text{ex}\mathcal{P}$. Indeed, it is immediate to notice that $\text{ex}\mathcal{P}=\{\mathcal{L}_{i}\}_{i=1}^{3}$. Let now $D_{N+1}\equiv D_{4}$ be a collection of $n$ samples from $P^{\text{true}}\equiv\mathcal{L}_{2}\in\mathcal{P}$. The hypotheses space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and we consider the zero-one loss $l$ as we did in the main portion of the paper.
@@
As we can see from Table B.1, our bound $\epsilon^{\star}(\delta)$ improves on the classical SLT one $\epsilon_{\text{UB}}(\delta)$. Table B.1 also tells us that bound $\epsilon^{\star}(\delta)$ is tighter than $\epsilon_{\text{UB}}(\delta)$ when the sample size $n=|D_{4}|$ is small, and then $\epsilon^{\star}(\delta)$ becomes progressively closer to $\epsilon_{\text{UB}}(\delta)$ as $n=|D_{4}|$ increases. This same pattern is observed when the extrema of the credal set are closer to each other. Indeed, in Table B.2 we repeat the experiment and choose as extrema of $\mathcal{P}$ three measures whose pdf’s are three Normals $\mathcal{N}(-0.1,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(0.1,1)$. [^10] Of course, they are much closer to each other than the Normals $\mathcal{N}(-5,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(5,1)$, e.g. in the Total Variation metric. The reason for this behavior is the following. With few available samples, that is, when $n=|D_{4}|$ is low, Credal Learning Theory is able to leverage the evidence encoded in the credal set, and hence to derive a tighter bound than classical Statistical Learning Theory. When the sample size is large, that is, when $n=|D_{4}|$ is high, the classical bound $\epsilon_{\text{UB}}(\delta)$ itself is very small. This is because the amount of evidence available is large, and so $1/n\sum_{i=1}^{n}l((x_{i},y_{i}),h)$ well approximates $\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\text{true}}(\text{d}(x,y))$. In turn, since $\epsilon_{\text{UB}}(\delta)$ is already very small, then the CLT bound $\epsilon^{\star}(\delta)$ we derive cannot improve greatly on it. Hence, their values are close together, despite $\epsilon^{\star}(\delta)$ being slightly tighter. The code for this experiment is available upon request.
@@
Table B.1: Results of experimental evaluation of our bound tightness. Here the hypotheses space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$.
@@
Table B.2: Results of experimental evaluation of our bound tightness. Here the hypotheses space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-0.1,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(0.1,1)$.
@@
For the experiment, we selected a standard Gaussian distribution $\mathcal{N}(0,1)$ (mean 00, standard deviation $1$) to generate data. Similarly to Experiment 1, (i) the hypotheses space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and (ii) we assume a zero-one loss function. The latter is used to evaluate the performance of the classifiers. For each run, we generated a training set and a test set from the standard Gaussian distribution. At training time, for each hypothesis $h$ in $\mathcal{H}$, we calculate the empirical risk on the training set using the zero-one loss and identify the hypothesis $\hat{h}$ that minimizes such risk (the empirical risk minimizer). At test time, we compute the empirical risk $L_{P}(\hat{h})$ of $\hat{h}$ on the test set as shown in Table B.3.111111Here $P$ denotes the probability measure whose pdf is the standard Normal density.
@@
Table B.3: Results of experimental evaluation. The hypotheses space is such that $|\mathcal{H}|=100$, and $\delta=0.05$.
We calculate the upper bound $\epsilon_{\text{UB}}(\delta)$ on the empirical risk based on Corollary 4.3, which is a function of the number of hypotheses in $\mathcal{H}$, the value of $\delta$, and the number $n$ of training samples. This is the classic SLT bound. The experiment is run $1000$ times with specified numbers of training and test samples, and a $\delta$ value of $0.05$ to check whether the condition (empirical risk on the test set is upper bounded by the theoretical bound) is satisfied in any of the $1000$ trials. The experimental results in Table B.3 validate the classical SLT bound by repeatedly testing it on randomly generated data. The code for this experiment is also available upon request.
Experiment 3: In this, we aim to empirically validate Theorem 4.5, which addresses the behavior of the empirical risk minimizer in the presence of a finite hypothesis space and no realizability. We generate synthetic data from Gaussian distributions (with the same parameters as in Experiment 1), with added uniform noise to ensure no realizability, meaning that no hypothesis can perfectly predict the labels. The hypotheses space $\mathcal{H}$ is defined as a set of threshold-based classifiers parameterized by $\theta$. For the experiment, we generate training samples $D_{1},D_{2},D_{3}$ and test sample $D_{4}$ from Gaussian distributions with added noise. Labels are created based on the samples, with noise introduced to flip labels randomly, ensuring that no hypothesis in $\mathcal{H}$ can achieve zero loss. We identify the empirical risk minimizer $\hat{h}$ using the combined training samples $D_{1},D_{2},D_{3}$. We calculate the empirical risk $L_{P}(\hat{h})$ using the test data $D_{4}$. The theoretical risk $L_{P}(h^{\star})$ is assumed to be the risk of a perfect classifier. We compute the theoretical bound $\epsilon^{\star\star}(\delta)$ and verify whether the difference $L_{P}(\hat{h})-L_{P}(h^{\star})$ is within this bound. The results show that the empirical risk of the empirical risk minimizer $\hat{h}$ is within the bound $\epsilon^{\star\star}(\delta)$ of the best theoretical model $h^{\star}$. The difference $L_{P}(\hat{h})-L_{P}(h^{\star})$ also satisfies the condition
$L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta).$
This validates empirically (in a synthetic environment) Theorem 4.5, showing that even under no realizability, the empirical risk minimizer’s performance is close to the theoretical best within a computable bound. The experimental results are presented in Table B.4, where we also show (i) that $\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)\doteq\sqrt{\frac{2\left(\log\left|\mathcal{H}\right|+\log\left(\frac{2}{\delta}\right)\right)}{n}}$ from Corollary 4.7 always holds; and (ii) that by foregoing realizability, we obtain a slightly looser bound. Indeed, as we can see, $\epsilon^{\star\star}(\delta)$ is slightly larger than $\epsilon^{\star}(\delta)$ from Table B.1 for all the sample size values $n$ that we consider. The code for this experiment is also available upon request.
# Samples $n$ $L_{P}(\hat{h})$ $L_{P}(h^{\star})$ $L_{P}(\hat{h})-L_{P}(h^{\star})$ $\epsilon^{\star\star}(\delta)$ $\epsilon^{\prime}_{\text{UB}}(\delta)$
$10$ $0.30000$ $0.10000$ $0.19999$ $0.76009$ $0.84877$
$100$ $0.14000$ $0.10000$ $0.04000$ $0.07600$ $0.26840$
$200$ $0.10500$ $0.10000$ $0.00499$ $0.03800$ $0.18979$
$300$ $0.11333$ $0.10000$ $0.01333$ $0.02533$ $0.15496$
$400$ $0.11750$ $0.10000$ $0.01749$ $0.01900$ $0.13420$
$500$ $0.10400$ $0.10000$ $0.00399$ $0.01520$ $0.12003$
Table B.4: Results of experimental evaluation of Theorem 4.5. Here the hypotheses space is such that $|\mathcal{H}|=100$, $\delta=0.05$ and noise level $0.1$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$. Let us remark that in this experiment we forego the assumption of realizability, that $L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)$ for every sample size we tested on, and that $\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)$ for every sample size we tested on.
@@
Then, suppose that $\epsilon_{1}=0.2$, $\epsilon_{2}=0.3$, $\epsilon_{3}=0.1$, and $\epsilon_{4}=0.25$, so that $\mathscr{L}_{1}=\{P:P=0.8\mathcal{L}_{1}+0.2Q\text{, }\forall Q\in\Delta_{\Omega}\}$, $\mathscr{L}_{2}=\{P:P=0.7\mathcal{L}_{2}+0.3Q\text{, }\forall Q\in\Delta_{\Omega}\}$, $\mathscr{L}_{3}=\{P:P=0.9\mathcal{L}_{3}+0.1Q\text{, }\forall Q\in\Delta_{\Omega}\}$, and $\mathscr{L}_{4}=\{P:P=0.75\mathcal{L}_{4}+0.25Q\text{, }\forall Q\in\Delta_{\Omega}\}$. By Wasserman and Kadane [74, Example 3], we know that, if we put $\mathcal{P}=\text{Conv}(\cup_{i=1}^{4}\mathscr{L}_{i})$, the following holds
$\underline{P}(A)=\begin{cases}\min_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A),&\forall A\neq\Omega\\ 1,&\text{if }A=\Omega\end{cases}$
and
$\overline{P}(A)=\begin{cases}\max_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)+\epsilon_{i},&\forall A\neq\emptyset\\ 0&\text{if }A=\emptyset\end{cases}.$